# SemTalk: Holistic Co-speech Motion Generation with Frame-level Semantic Emphasis

### v1

@@ 5 Conclusion
We propose SemTalk, a novel approach for holistic co-speech motion generation with frame-level semantic emphasis. Our method addresses the integration of sparse yet crucial semantically significant gestures into foundational rhythm-related motions, which has received less attention in previous works. We develop a framework that separately learns rhythm-related base motion through rhythmic consistency learning and semantic-aware semantic motion through semantic emphasis learning, and adaptively fuses them based on a learned semantic score. Our approach has demonstrated state-of-the-art performance on two public datasets quantitatively and qualitatively. The qualitative results and user study show that our method can generate high-quality co-speech motion sequences that enhance frame-level semantics over robust base motions, reflecting the full spectrum of human expressiveness.

### v3

@@ 5 Conclusion
We propose SemTalk, a novel approach for holistic co-speech motion generation with frame-level semantic emphasis. Our method addresses the integration of sparse yet expressive motion into foundational rhythm-related motion, which has received less attention in previous works. We develop a framework that separately learns rhythm-related base motion through coarse2fine cross-attention module and rhythmic consistency learning, while capturing semantic-aware motion through Semantic Emphasis Learning. These components are then adaptively fused based on a learned semantic score. Our approach has demonstrated state-of-the-art performance on two public datasets quantitatively and qualitatively. The qualitative results and user study show that our method can generate high-quality co-speech motion sequences that enhance frame-level semantics over robust base motions, reflecting the full spectrum of human expressiveness.